# On some extensions of shape-constrained generalized additive modelling in R

### 14 Mar 2024 v1

@@ 3.1 SCAM with smooth interactions
Here, $f_{1}$ and $f_{2}$ represent smooth ‘main effects’. $m_{3}$ is a smooth ‘interaction’ subject to, for example, an increasing constraint with respect to $x$ and the main effects excluded ($f_{1}$ can also be subject to an increasing constraint). The construction of such a functional ANOVA decomposition of shape-constrained smooths relies on constructing a tensor product basis. This recognizes the equivalence between tensor product basis construction and the method used for interactions in linear models. Regression modelling with interactions within generalized additive models, where the will is to separate the unconstrained smooth interaction effect of two or more covariates from the main smooth effects (ANOVA decompositions of smooths), was introduced in Wood [2017]. Following that work, the marginal smooths of a tensor product are subject to identifiability constraints before constructing the tensor product basis. This leads to interaction smooths that do not encompass the corresponding main effects. Two tensor product bivariate interaction smooths under shape constraints have been developed and implemented in the R package scam, with increasing and decreasing constraints in the first covariates. Both smooth interactions apply the corresponding SCOP identifiability constraints to the first marginal and centering constraints to the second unconstrained marginal. The SCOP identifiability and centering constraints remove the constant functions from the bases of the marginals. By doing so, constant functions are eliminated from the marginal bases, preventing the interaction term from containing the main effects of $f_{1}$ and $f_{2}$. This would happen if the marginal bases were multiplied by the constant functions in the other marginal bases. The removal of constraints does not impact penalty terms, given that constant functions lie within the null space of the penalty. In the scam package, two shape-constrained smooth interaction terms with increasing and decreasing constraints are implemented.
@@ 3.4 Smoothing parameter estimation
The current fitting approach for SCAM is reasonably efficient and robust when applied to data sets containing a few tens of thousands of observations. If $n$ and $p$ denote numbers of rows and columns of the model matrix correspondingly, then the approach has the $O(np^{2})$ computational cost of standard penalized regression spline based GAM estimation and typically involves 2-4 times as many $O(np^{2})$ steps due to the additional non-linearities required for dealing with the shape constrained terms. As an alternative to the conventional quasi-Newton approach for smoothing parameter $\bm{\lambda}$ optimization, an extended Fellner-Schall method (EFS) reported in Wood and Fasiolo [2017] was introduced. The smoothing parameters are estimated in the outer optimization scheme using a full Newton method for coefficient estimation. The extended Fellner-Schall method generalizes the technique of Fellner [1986] and Schall [1991] and offers a significant simplification over existing methods by re-using quantities anyway required for model coefficients estimation when iteratively updating $\bm{\lambda}$. When modelling with SCAM, rather than maximizing the log restricted marginal likelihood as it is proposed in [Wood and Fasiolo, 2017], similar convenient updates are obtained for minimizing a prediction error GCV criterion, $\mathcal{V}_{g}=nD(\hat{\bm{\beta}})/(n-\gamma\tau)^{2},$ when the scale parameter of the exponential family distribution is unknown and for UBRE, $\mathcal{V}_{a}=D(\hat{\bm{\beta}})+2\phi\tau\gamma,$ for known scale parameter. The smoothing parameter update for GCV minimization is
$\lambda_{j}^{*}=-\frac{2\gamma D(\bm{\beta})}{(n-\tau)}\cdot\frac{\partial\tau}{\partial\lambda_{j}}\cdot\left(\frac{\partial D(\bm{\beta})}{\bm{\beta}}\frac{d\hat{\bm{\beta}}}{d\lambda_{j}}\right)^{-1}\cdot\lambda_{j},$
where $D$ is the model deviance, $\tau$ is the effective degrees of freedom, $\gamma$ is a parameter that usually has the value $1,$ but can be increased to get smoother models. The $\lambda$ update when minimizing UBRE is
$\lambda_{j}^{*}=-2\phi\gamma\cdot\frac{\partial\tau}{\partial\lambda_{j}}\cdot\left[\left(\frac{\partial D}{\partial\bm{\beta}}\right)^{T}\cdot\frac{d\hat{\bm{\beta}}}{d\lambda_{j}}\right]^{-1}\cdot\lambda_{j}$
This leads to a more robust smoothing parameter scheme for SCAM that eliminates the need for implicit differentiation in quasi-Newton optimization. The EFS approach is implemented in the scam package and is called for by optimizer="efs" argument of the scam() function.

### 17 Mar 2024 v2

@@ 3.1 SCAM with smooth interactions
Here, $f_{1}$ and $f_{2}$ represent smooth ‘main effects’. $m_{3}$ is a smooth ‘interaction’ subject to, for example, an increasing constraint with respect to $x$ and the main effects excluded ($f_{1}$ can also be subject to an increasing constraint). The construction of such a functional ANOVA decomposition of shape-constrained smooths relies on constructing a tensor product basis. This recognizes the equivalence between tensor product basis construction and the method used for interactions in linear models. Regression modelling with interactions within generalized additive models, where the will is to separate the unconstrained smooth interaction effect of two or more covariates from the main smooth effects (ANOVA decompositions of smooths), was introduced in Wood [2017]. Following that work, the marginal smooths of a tensor product are subject to identifiability constraints before constructing the tensor product basis. This leads to interaction smooths that do not encompass the corresponding main effects.
Two tensor product bivariate interaction smooths under shape constraints have been developed and implemented in the R package scam, with increasing and decreasing constraints in the first covariates. Both smooth interactions apply the corresponding SCOP identifiability constraints to the first marginal and centering constraints to the second unconstrained marginal. The SCOP identifiability and centering constraints remove the constant functions from the bases of the marginals. By doing so, constant functions are eliminated from the marginal bases, preventing the interaction term from containing the main effects of $f_{1}$ and $f_{2}$. This would happen if the marginal bases were multiplied by the constant functions in the other marginal bases. The removal of constraints does not impact penalty terms, given that constant functions lie within the null space of the penalty. In the scam package, two shape-constrained smooth interaction terms with increasing and decreasing constraints are implemented.
@@ 3.4 Smoothing parameter estimation
The current fitting approach for SCAM is reasonably efficient and robust when applied to data sets containing a few tens of thousands of observations. If $n$ and $p$ denote numbers of rows and columns of the model matrix correspondingly, then the approach has the $O(np^{2})$ computational cost of standard penalized regression spline based GAM estimation and typically involves 2-4 times as many $O(np^{2})$ steps due to the additional non-linearities required for dealing with the shape constrained terms. As an alternative to the conventional quasi-Newton approach for smoothing parameter $\bm{\lambda}$ optimization, an extended Fellner-Schall method (EFS) reported in Wood and Fasiolo [2017] was introduced. The smoothing parameters are estimated in the outer optimization scheme using a full Newton method for coefficient estimation. The extended Fellner-Schall method generalizes the technique of Fellner [1986] and Schall [1991] and offers a significant simplification over existing methods by re-using quantities anyway required for model coefficients estimation when iteratively updating $\bm{\lambda}$. When modelling with SCAM, rather than maximizing the log restricted marginal likelihood as it is proposed in [Wood and Fasiolo, 2017], similar convenient updates are obtained for minimizing a prediction error GCV criterion, $\mathcal{V}_{g}=nD(\hat{\bm{\beta}})/(n-\gamma\tau)^{2},$ when the scale parameter of the exponential family distribution is unknown and for the UBRE, $\mathcal{V}_{a}=D(\hat{\bm{\beta}})+2\phi\tau\gamma,$ for known scale parameter. The smoothing parameter update for the GCV minimization is
$\rho_{j}^{*}=\rho_{j}+\log\left\{-\frac{2\gamma D}{n-\tau}\cdot\frac{\partial\tau}{\partial\rho_{j}}\left[\left(\frac{\partial D}{\partial\bm{\beta}}\right)^{T}\cdot\frac{d\hat{\bm{\beta}}}{d\rho_{j}}\right]^{-1}\right\},$
where $\rho_{j}=\log(\lambda_{j})$ to ensure that $\lambda_{j}^{*}$ is positive, $D$ is the model deviance, $\tau$ is the effective degrees of freedom, $\gamma$ is a parameter that usually has the value $1,$ but can be increased to get smoother models. The $\rho$ update when minimizing the UBRE score is
$\rho_{j}^{*}=\rho_{j}+\log\left\{-2\phi\gamma\cdot\frac{\partial\tau}{\partial\rho_{j}}\left[\left(\frac{\partial D}{\partial\bm{\beta}}\right)^{T}\cdot\frac{d\hat{\bm{\beta}}}{d\rho_{j}}\right]^{-1}\right\}.$
The EFS approach eliminates the need for implicit differentiation in quasi-Newton optimization and leads to a more robust smoothing parameter scheme for SCAM. It is implemented in the scam package and is called for by optimizer="efs" argument of the scam() function.